7. A drawing pin is pushed against a wooden table with a force of 10N. Calculate the
pressure exerted by the pin at a point on a table, if the area of the point is 180cm2.

Grade:8

1 Answers

Tanu Kajla
2 years ago
Force=10N
Area=180cm2.    Convert it into m=180/10000=0.0180m2
Pressure=force/area=10/0.0180=555.5pascal